# Improving variational counterdiabatic driving with weighted actions and computer algebra

### v1

@@ Appendix A Derivations of equations
Partial action decomposition
We decompose the action into the partial action in Eq. (32). By combining the expansion of the action in Eq. (14) and the definition of the weight in Eq. (15), we obtain
$\mathcal{S}^{(K)}[V]=\frac{1}{\hbar^{2}}\sum_{nm}(\epsilon_{m}-\epsilon_{n})^{2}w_{mn}^{(K)}\bigl{|}[V-\Phi(\lambda)]_{mn}\bigr{|}^{2}+\mathrm{const.}$ (48)
The summand is symmetric under the exchange of $m$ and $n$, and it is zero for $m=n$. Therefore, we can combine the $(n,m)$ and $(m,n)$ terms to rewrite
$\mathcal{S}^{(K)}[V]=\frac{2}{\hbar^{2}}\sum_{\begin{subarray}{c}nm\\ w_{m}^{(K)}\leq w_{n}^{(K)}\end{subarray}}(\epsilon_{m}-\epsilon_{n})^{2}\chi_{mn}^{(K)}w_{mn}^{(K)}\bigl{|}[V-\Phi(\lambda)]_{mn}\bigr{|}^{2}$
$\hskip 150.00023pt+\mathrm{const.},$ (49)
where we insert $2\chi_{mn}^{(K)}$ because if $m$ and $n$ satisfy $w_{m}^{(K)}=w_{n}^{(K)}$ and $m\neq n$, both $(m,n)$ and $(n,m)$ appear in the sum $\sum_{nm:w_{m}^{(K)}\leq w_{n}^{(K)}}$. The decomposition in Eq. (32) easily follows from Eq. (49).
Next, we show that the $K$-dependence of the partial action $\mathcal{T}^{(K,n)}[V]$ is between $O(K^{0})$ and $O(K^{-2})$. For simplicity, we neglect the $K$-dependence of $E^{(K)}$ because the dependence is subtle [see, e.g., Fig. 1(a)]. Then, the $K$-dependence of $\mathcal{T}^{(K,n)}[V]$ is mainly captured by the ratio $w_{mn}^{(K)}/w_{n}^{(K)}$. This ratio is written as
$\frac{w_{mn}^{(K)}}{w_{n}^{(K)}}=\left[\frac{1}{K}\sum_{s=0}^{K-1}\left(\frac{\epsilon_{m}-E^{(K)}}{\epsilon_{n}-E^{(K)}}\right)^{s}\right]^{2},$ (50)
where we used the second line of Eq. (43). The condition $w_{m}^{(K)}\leq w_{n}^{(K)}$ implies $|\epsilon_{m}-E^{(K)}|\leq|\epsilon_{n}-E^{(K)}|$, and thus, the sum in Eq. (50) is a finite geometric sum with a common ratio $(\epsilon_{m}-E^{(K)})/(\epsilon_{n}-E^{(K)})\in[-1,1]$. From Eq. (50), we find the scaling of the ratio $w_{mn}^{(K)}/w_{n}^{(K)}$ in three extreme cases: (i) if $|\epsilon_{m}-E^{(K)}|\ll|\epsilon_{n}-E^{(K)}|$, the geometric sum converges quickly for a small value of $K$, and thus the ratio scales as $O(K^{-2})$; (ii) if $(\epsilon_{m}-E^{(K)})\simeq(\epsilon_{n}-E^{(K)})$ and $K$ is not too large, the geometric sum is roughly proportional to $K$, and the ratio scales as $O(K^{0})$; However, as $K$ becomes larger, the geometric sum starts to converge, and the scale changes to $O(K^{-2})$; (iii) if $(\epsilon_{m}-E^{(K)})\simeq-(\epsilon_{n}-E^{(K)})$, the sign alternates in the geometric sum, and the value of $w_{mn}^{(K)}/w_{n}^{(K)}$ oscillates for a small $K$; Nevertheless, the scaling eventually becomes $O(K^{-2})$. A more general case lies between these extreme cases. The scaling of $\mathcal{T}^{(K,n)}[V]$ is determined by the sum of many terms with these scalings, and thus, the overall scaling also lies between $O(K^{0})$ and $O(K^{-2})$.
The asymptotic formula of $\mathcal{T}^{(K,n)}[V]$ is explicitly obtained from the first line of Eq. (43). By using $(\epsilon_{n}-E^{(K)})^{K}\gg(\epsilon_{m}-E^{(K)})^{K}$ for a sufficiently large $K$ in the first line of Eq. (43), the ratio $w_{mn}^{(K)}/w_{n}^{(K)}$ is approximated as
$\frac{w_{mn}^{(K)}}{w_{n}^{(K)}}\simeq\frac{1}{K^{2}}\frac{(\epsilon_{n}-E^{(K)})^{2}}{(\epsilon_{m}-\epsilon_{n})^{2}}.$ (51)
Using this approximation, the partial action $\mathcal{T}^{(K,1)}[V]$ is approximated by $K^{-2}(\epsilon_{1}-E^{(K)})^{2}\sum_{m:\,m\neq 1}|[V-\Phi]_{1m}|^{2}$, where we used $\chi_{1m}^{(K)}=1$ for a sufficiently large $K$ because $w_{1}^{(K)}$ becomes larger than any other $w_{m}^{(K)}$. This asymptotic formula of $\mathcal{T}^{(K,1)}[V]$ has been used in the main text to show that $\mathcal{S}^{(K)}[V]$ approaches the ideal action up to normalization.
The ideal action can also be derived from the algebraic expression of $\mathcal{S}^{(K)}[V]$ in Eq. (22a). In the limit $K\to\infty$, the power of the shifted Hamiltonian,
$(H-E^{(K)})^{K}=\sum_{n=1}^{D}(\epsilon_{n}-E^{(K)})^{K}|\phi_{n}\rangle\langle\phi_{n}|,$ (52)
converges to the projection operator to the ground state $|\phi_{1}\rangle\langle\phi_{1}|$ up to an overall coefficient. Therefore, the action $\mathcal{S}^{(K)}[V]$ converges to
$\bigl{\|}\partial_{\lambda}(|\phi_{1}\rangle\langle\phi_{1}|)-i\hbar^{-1}[|\phi_{1}\rangle\langle\phi_{1}|,V]\bigr{\|}^{2}=\frac{2}{\hbar^{2}}\sum_{n\neq 1}\bigl{|}[V-\Phi]_{1n}\bigr{|}^{2}$ (53)
up to normalization, which reproduces the ideal action in Eq. (33).
Appendix B Appendix to the general framework in Sec. II
Locality of the conventional action
We show that the conventional action is given by a sum of local terms, assuming that $H(\lambda)$, $\partial_{\lambda}H(\lambda)$, and $A_{\mu}$’s are $k$-local operators with $k$ small and independent of the system size. More precisely, for every fixed $\lambda$, we assume that we can write $H(\lambda)=h_{1}+h_{2}+\cdots$, $\partial_{\lambda}H(\lambda)=l_{1}+l_{2}+\cdots$, and $A_{\mu}=a_{\mu 1}+a_{\mu 2}+\cdots$, where $h_{i}$, $l_{i}$, and $a_{\mu i}$ act nontrivially on at most $k$ sites. Based on this assumption, we can expand the action $\mathcal{S}_{\lambda}^{(1)}[V]$ in Eq. (6b) as
$\mathcal{S}_{\lambda}^{(1)}[V]$ $=-\hbar^{-2}\sum_{\mu ij}\sum_{\nu i^{\prime}j^{\prime}}\mathop{\mathrm{tr}}\left\{[h_{i},a_{\mu j}][h_{i^{\prime}},a_{\nu j^{\prime}}]\right\}\alpha_{\mu}\alpha_{\nu}$
In this expansion, $[h_{i},a_{\mu j}]$ is zero if the domain of $h_{i}$ does not overlap the domain of $a_{\mu j}$. Otherwise, $[h_{i},a_{\mu j}]$ is an at most $(2k-1)$-local operator. Furthermore, if the domains of $[h_{i},a_{\mu j}]$ and $l_{i^{\prime}}$ do not overlap, the trace $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}]l_{i^{\prime}}\}$ is proportional to the product $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}]\}\mathop{\mathrm{tr}}(l_{i^{\prime}})$, which is zero due to $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}]\}=0$. Otherwise, the operator $[h_{i},a_{\mu j}]l_{i^{\prime}}$ is at most $(3k-2)$-local, and thus the trace $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}]l_{i^{\prime}}\}$ is determined by at most $(3k-2)$-local information. Similarly, the trace $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}][h_{i^{\prime}},a_{\nu j^{\prime}}]\}$ is either zero or dependent on at most $(4k-3)$-local information. Thus, the action $\mathcal{S}_{\lambda}^{(1)}[V]$ consists of local terms except for the “const.” term, which does not affect the minimization.
This locality of $\mathcal{S}_{\lambda}^{(1)}[V]$ does not necessarily imply that the resulting coefficient $\alpha_{\mu}$ depends only on local information. Nevertheless, this is true if $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}][h_{i^{\prime}},a_{\nu j^{\prime}}]\}=0$ for any $\mu\neq\nu$, in which case the determination of $\alpha_{\mu}$ is decoupled from the determination of $\alpha_{\nu}$. This situation arises, for example, for locally interacting spin systems with the one-body driving [27, 28].
Use of a conserved charge
When the Hamiltonian has a conserved charge, our formulation can incorporate it to derive a wider variety of weighted actions. Let $J(\lambda)$ be an operator that satisfies $[H(\lambda),J(\lambda)]=0$, and we expand it in the Hamiltonian basis as
Using a bivariate degree-$K$ polynomial $\mathcal{P}_{\lambda}(x,y)=\sum_{k=0}^{K}\sum_{l=0}^{k}p_{l,k-l}(\lambda)x^{l}y^{k-l}$, we can use a weighted action of the form,
×$\qquad\qquad\qquad\times\bigl{|}[V-\Phi(\lambda)]_{mn}\bigr{|}^{2}\Bigr{\}}+\mathrm{const.}$ (56)
For example, let us consider that $J(\lambda)$ is a parity operator independent of $\lambda$ with the eigenvalues $j_{n}(\lambda)=\pm 1$, and thus the eigenstates are split into even- and odd-parity states. Let us also assume that the driving operator $A_{\mu}$ has the even parity for all $\mu$, meaning that $JA_{\mu}J=A_{\mu}$. Then, the matrix element $[V-\Phi(\lambda)]_{mn}$ is nonzero only when the states $m$ and $n$ have the same parity. In this case, we can use $\mathcal{P}_{\lambda}(H(\lambda),J)=[H(\lambda)-E^{(K)}(\lambda)]^{K}(J+1)$ to prioritize the suppression of nonadiabatic transitions between even-parity states. With this choice, $\mathcal{P}_{\lambda}(\epsilon_{n}(\lambda),j_{n}(\lambda))$ is zero for any odd-parity $n$, and thus the weighted action only penalizes the nonadiabatic transitions between even-parity states.
Efficient algorithm for the entire range of $\lambda$
We discuss an efficient algorithm to compute the driving protocol over the whole range of $\lambda$, assuming that the Hamiltonian has the form of Eq. (19). We first introduce some notation. Let $\tilde{f}_{1}(\lambda),\tilde{f}_{2}(\lambda),\dots$ be the list of all the possible $k$-fold products of the coefficients $\{f_{1}(\lambda),\dots,f_{\Gamma}(\lambda)\}$ for $k=0,\dots,K$. For example, when $K=2$ and $\Gamma=2$, the list is $(\tilde{f}_{1}(\lambda),\dots,\tilde{f}_{6}(\lambda))=(f_{1}(\lambda)^{2},f_{1}(\lambda)f_{2}(\lambda),f_{2}(\lambda)^{2},f_{1}(\lambda),f_{2}(\lambda),1)$. The order of the list does not matter. We define $\nu_{g}$ as the total number of $f_{\gamma}$’s contained in $\tilde{f}_{g}(\lambda)$. In the example above, the numbers are $(\nu_{1},\dots,\nu_{6})=(2,2,2,1,1,0)$. We then define $\lambda$-independent operators $\tilde{F}_{1},\tilde{F}_{2},\dots$ by the relation,
$H(\lambda)^{k}=\left(\sum_{\gamma=1}^{\Gamma}f_{\gamma}(\lambda)F_{\gamma}\right)^{k}\eqqcolon\sum_{g:\,\nu_{g}=k}\tilde{f}_{g}(\lambda)\tilde{F}_{g}$ (57)
for $k=0,\dots,K$. In other words, $\tilde{F}_{g}$ is defined as the terms in $H(\lambda)^{k}$ whose $\lambda$-dependence is given by $\tilde{f}_{g}(\lambda)$. In the example above, the operators are $(\tilde{F}_{1},\dots,\tilde{F}_{6})=((F_{1})^{2},F_{1}F_{2}+F_{2}F_{1},(F_{2})^{2},F_{1},F_{2},I)$, where $I$ is the identity.
Using these definitions, we can conveniently separate $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ into $\lambda$-dependent coefficients and traces of $\lambda$-independent operators. By inserting Eqs. (10) and (57) into the definitions of $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ in Eq. (17), we get
$Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ $=-\sum_{gg^{\prime}}p_{\nu_{g}}(\lambda)p_{\nu_{g^{\prime}}}(\lambda)\tilde{f}_{g}(\lambda)\tilde{f}_{g^{\prime}}(\lambda)\tilde{Q}_{\mu\nu,gg^{\prime}},$ (58a)
$r_{\mu}^{(\mathcal{P})}(\lambda)$ $=i\hbar\sum_{gg^{\prime}}p_{\nu_{g}}(\lambda)p_{\nu_{g^{\prime}}}(\lambda)[\partial_{\lambda}\tilde{f}_{g}(\lambda)]\tilde{f}_{g^{\prime}}(\lambda)\tilde{r}_{\mu\nu,gg^{\prime}},$ (58b)
where we define
$\tilde{r}_{\mu\nu,gg^{\prime}}$ $\coloneqq\mathop{\mathrm{tr}}\left\{\tilde{F}_{g}[\tilde{F}_{g^{\prime}},A_{\mu}]\right\}.$ (59b)
These expressions allow efficient computation of $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$. To calculate them over the entire range of $\lambda$, we only need to calculate the traces $\tilde{Q}_{\mu\nu,gg^{\prime}}$ and $\tilde{r}_{\mu\nu,gg^{\prime}}$ once.
Appendix C Appendix to the ground-state evolution in Sec. III
@@
Figure 8: Typical behavior of the function $\Omega_{\lambda}^{(K)}(E)$ and the weight $[\epsilon_{n}(\lambda)-E]^{2K-2}$, exemplified with a ferromagnetic system of size $N=12$ with $K=3$ at $\lambda=0.25$. (a) Typical shape of the function $\Omega_{\lambda}^{(K)}(E)$ (red), which has one maximum and one minimum. The optimal energy shift $E_{\lambda}^{(K)}$ is determined as the value of $E$ that minimizes $\Omega_{\lambda}^{(K)}(E)$ (black dashed line). The energies $\epsilon_{1}$ and $\epsilon_{D}$ denote the lowest and highest energy eigenvalues. (b) To understand this shape of $\Omega_{\lambda}^{(K)}(E)$, we plot the weight $[\epsilon_{n}(\lambda)-E]^{2K-2}$ for six different values of $E$ (green–blue). Green corresponds to a smaller $E$, and blue corresponds to a larger $E$, as indicated by the arrows in panel (a). Each weight is normalized so that the area under the curve is constant for better visualization, noting that the overall normalization is unimportant. Gray histogram shows the density of states.
@@
In Fig. 8(a), we show a typical shape of the function $\Omega_{\lambda}^{(K)}(E)$, which has one maximum and one minimum. To understand this shape, we recall the definition of $\Omega_{\lambda}^{(K)}(E)$ in Eq. (25b), which says that $\Omega_{\lambda}^{(K)}(E)$ is the weighted average of the energy eigenvalues $\epsilon_{n}(\lambda)$ over the weight $[\epsilon_{n}(\lambda)-E]^{2K-2}$. We plot the weight for six values of $E$ in Fig. 8(b). When $E$ takes a large negative value ($\epsilon_{1}-E\gg\epsilon_{D}-\epsilon_{1}$), the weight is approximated as $[\epsilon_{n}(\lambda)-E]^{2K-2}\simeq E^{2K-2}$, and thus, it is almost uniform over the entire range of the energy spectrum. As $E$ is gradually increased, the weight becomes an increasing function of the energy eigenvalue over the energy spectrum, putting weight to the high-energy states [green curves in Fig. 8(b)]. As $E$ is further increased, the minimum point of the weight passes through the middle of the energy spectrum. Then, the weight turns into a decreasing function of the energy eigenvalue over the energy spectrum, and the weight concentrates on the low-energy states [blue curves in Fig. 8(b)]. Finally, as $E$ becomes larger ($E-\epsilon_{D}\gg\epsilon_{D}-\epsilon_{1}$), the weight becomes almost uniform again. This course of change forms the maximum and minimum of $\Omega_{\lambda}^{(K)}(E)$.
The optimal energy shift $E_{\lambda}^{(K)}$ is determined so that the weight $[\epsilon_{n}(\lambda)-E]^{2K-2}$ is most concentrated on the low-energy states. Due to the shape of $\Omega_{\lambda}^{(K)}(E)$, the resulting $E_{\lambda}^{(K)}$ is typically higher than the center of the energy spectrum. This behavior is confirmed by Fig. 1(a).
The numerical minimization of $\Omega_{\lambda}^{(K)}(E)$ is a one-dimensional optimization problem, and it can be stably solved by the following procedure. We start with Eq. (26), which shows that $\Omega_{\lambda}^{(K)}(E)$ is a rational function of $E$. We analytically calculate $d\Omega_{\lambda}^{(K)}/dE$, which will again be a rational function of $E$, and equate the numerator to zero. The numerator is a polynomial, and the resulting polynomial equation is expected to have two real roots because $\Omega_{\lambda}^{(K)}(E)$ has two extremum points. We need to solve the larger root of the polynomial equation. For $K=2$, the polynomial equation is quadratic and analytically solvable, and the larger solution gives $E^{(2)}_{\lambda}$. For $K\geq 3$, we can use standard numerical methods, such as the Newton method, to get $E_{\lambda}^{(K)}$. As the initial value of the Newton method, we can use the analytical value of $E^{(2)}_{\lambda}$, which is expected to be close to $E_{\lambda}^{(K)}$ with $K\geq 3$.
@@
When the Hamiltonian has the form of Eq. (19), we can efficiently evaluate $\omega_{k}(\lambda)=\mathop{\mathrm{tr}}[H(\lambda){}^{k}]$ for $k=1,\dots,2K-1$ by separating the $\lambda$-dependent coefficients and $\lambda$-independent operators, as has been done for $Q_{\mu\nu}^{(K)}(\lambda)$ and $r_{\mu}^{(K)}(\lambda)$ in the general framework in Appendix B.3. We expand $\omega_{k}(\lambda)=\mathop{\mathrm{tr}}[H(\lambda)^{k}]$ as follows: For $k=1,\dots,K$,
@@
$\omega_{k}(\lambda)={\displaystyle\sum_{g:\,\nu_{g}=k}}\tilde{f}_{g}(\lambda)\tilde{\omega}_{g},\qquad\tilde{\omega}_{g}\coloneqq\mathop{\mathrm{tr}}(\tilde{F}_{g}),$ (60)
and for $k=K+1,\dots,2K-1$,
$\omega_{k}(\lambda)$ $={\displaystyle\sum_{g:\,\nu_{g}=K}\,\sum_{g^{\prime}:\,\nu_{g^{\prime}}=k-K}}\!\!\tilde{f}_{g}(\lambda)\tilde{f}_{g^{\prime}}(\lambda)\tilde{\omega}_{gg^{\prime}},$ (61a)
$\tilde{\omega}_{gg^{\prime}}\,$ $\!\coloneqq\mathop{\mathrm{tr}}(\tilde{F}_{g}\tilde{F}_{g^{\prime}}).$ (61b)
@@
Once we compute the traces $\tilde{\omega}_{g}$ and $\tilde{\omega}_{gg^{\prime}}$, we can easily calculate $\omega_{k}(\lambda)$ over the entire range of $\lambda$.
Appendix D Details of the computer algebra for general spin-1/2 systems

### v3

@@ Appendix B Appendix to the ground-state evolution
Figure 9: Typical behavior of the function $\Omega_{\lambda}^{(K)}(E)$ and the weight $\epsilon_{n}(\lambda)-E]^{2K-2}$, exemplified with a ferromagnetic system of size $N=12$ with $K=3$ at $\lambda=0.25$. (a) Typical shape of the function $\Omega_{\lambda}^{(K)}(E)$ (red), which has one maximum and one minimum. The optimal energy shift $E_{\lambda}^{(K)}$ is determined as the value of $E$ that minimizes $\Omega_{\lambda}^{(K)}(E)$ (black dashed line). The energies $\epsilon_{1}$ and $\epsilon_{D}$ denote the lowest and highest energy eigenvalues. (b) To understand this shape of $\Omega_{\lambda}^{(K)}(E)$, we plot the weight $\epsilon_{n}(\lambda)-E]^{2K-2}$ for six different values of $E$ (green–blue). Green corresponds to a smaller $E$, and blue corresponds to a larger $E$, as indicated by the arrows in panel (a). Each weight is normalized so that the area under the curve is constant for better visualization, noting that the overall normalization is unimportant. Gray histogram shows the density of states.
@@
In Fig. 9(a), we show a typical shape of the function $\Omega_{\lambda}^{(K)}(E)$, which has one maximum and one minimum. To understand this shape, we recall the definition of $\Omega_{\lambda}^{(K)}(E)$ in Eq. (24b), which says that $\Omega_{\lambda}^{(K)}(E)$ is the weighted average of the energy eigenvalues $\epsilon_{n}(\lambda)$ over the weight $\epsilon_{n}(\lambda)-E]^{2K-2}$. We plot the weight for six values of $E$ in Fig. 9(b). When $E$ takes a large negative value ($\epsilon_{1}-E\gg\epsilon_{D}-\epsilon_{1}$), the weight is approximated as $\epsilon_{n}(\lambda)-E]^{2K-2}\simeq E^{2K-2}$, and thus, it is almost uniform over the entire range of the energy spectrum. As $E$ is gradually increased, the weight becomes an increasing function of the energy eigenvalue over the energy spectrum, putting weight to the high-energy states [green curves in Fig. 9(b)]. As $E$ is further increased, the minimum point of the weight passes through the middle of the energy spectrum. Then, the weight turns into a decreasing function of the energy eigenvalue over the energy spectrum, and the weight concentrates on the low-energy states [blue curves in Fig. 9(b)]. Finally, as $E$ becomes larger ($E-\epsilon_{D}\gg\epsilon_{D}-\epsilon_{1}$), the weight becomes almost uniform again. This course of change forms the maximum and minimum of $\Omega_{\lambda}^{(K)}(E)$.
The optimal energy shift $E=E_{\lambda}^{(K)}$ is determined so that the weight $\epsilon_{n}(\lambda)-E]^{2K-2}$ is the most concentrated on the low-energy states. Due to the shape of $\Omega_{\lambda}^{(K)}(E)$, the resulting $E_{\lambda}^{(K)}$ is typically higher than the center of the energy spectrum. This behavior is confirmed by Fig. 2(a).
The numerical minimization of $\Omega_{\lambda}^{(K)}(E)$ is a one-dimensional optimization problem, and it can be stably solved by the following procedure. We start with Eq. (25), which shows that $\Omega_{\lambda}^{(K)}(E)$ is a rational function of $E$. We analytically calculate $d\Omega_{\lambda}^{(K)}/dE$, which will again be a rational function of $E$, and we equate the numerator to zero. The numerator is a polynomial of $E$, and the resulting polynomial equation is expected to have two real roots because $\Omega_{\lambda}^{(K)}(E)$ has two extremum points. We need to solve the larger root of the polynomial equation. For $K=2$, the polynomial equation is quadratic and analytically solvable, and the larger solution gives $E^{(2)}_{\lambda}$. For $K\geq 3$, we can use standard numerical methods, such as the Newton method, to get $E_{\lambda}^{(K)}$. As the initial value of the Newton method, we can use the analytical value of $E^{(2)}_{\lambda}$, which is expected to be close to $E_{\lambda}^{(K)}$ with $K\geq 3$.
@@
When the Hamiltonian has the form of Eq. (19), we can efficiently evaluate $\omega_{k}(\lambda)=\mathop{\mathrm{tr}}[H(\lambda){}^{k}$] for $k=1,\dots,2K-1$ by separating the $\lambda$-dependent coefficients and $\lambda$-independent operators, as has been done for $Q_{\mu\nu}^{(K)}(\lambda)$ and $r_{\mu}^{(K)}(\lambda)$ in the general framework in Appendix A.4. We use Eq. (45) to expand $\omega_{k}(\lambda)=\mathop{\mathrm{tr}}[H(\lambda)^{k}]$. For $k=1,\dots,K$, we obtain
@@
$\omega_{k}(\lambda)={\displaystyle\sum_{g:\,d_{g}=k}}\tilde{f}_{g}(\lambda)\tilde{\omega}_{g},\qquad\tilde{\omega}_{g}\coloneqq\mathop{\mathrm{tr}}(\tilde{F}_{g}),$ (53)
where $\tilde{f}_{g}(\lambda)$ and $\tilde{F}_{g}$ are the scalar coefficient and operator, respectively, defined by the expansion of $H(\lambda)^{k}$ in Eq. (45). Similarly, for $k=K+1,\dots,2K-1$, we obtain
$\omega_{k}(\lambda)$ $={\displaystyle\sum_{g:\,d_{g}=K}\,\sum_{g^{\prime}:\,d_{g^{\prime}}=k-K}}\!\!\tilde{f}_{g}(\lambda)\tilde{f}_{g^{\prime}}(\lambda)\tilde{\omega}_{gg^{\prime}},$ (54a)
$\tilde{\omega}_{gg^{\prime}}\,$ $\!\coloneqq\mathop{\mathrm{tr}}(\tilde{F}_{g}\tilde{F}_{g^{\prime}}),$ (54b)
@@
where we used $\omega_{k}(\lambda)=\mathop{\mathrm{tr}}[H(\lambda)^{K}H(\lambda)^{k-K}]$. Once we compute the traces $\tilde{\omega}_{g}$ and $\tilde{\omega}_{gg^{\prime}}$, we can easily calculate $\omega_{k}(\lambda)$ over the entire range of $\lambda$.
Derivation of the partial action decomposition and the ideal action
We derive the partial action decomposition and its properties used in Sec. V.1. We fix $\lambda$ and omit the $\lambda$ dependencies of quantities.
To decompose the action into partial actions in Eq. (31), we combine the expansion of the action in Eq. (14) and the definition of the weight in Eq. (15) to obtain
$\mathcal{S}^{(K)}[V]=\frac{1}{\hbar^{2}}\sum_{nm}(\epsilon_{m}-\epsilon_{n})^{2}w_{mn}^{(K)}\bigl|[V-\Phi]_{mn}\bigr|^{2}+\mathrm{const.}$ (55)
The summand is symmetric under the exchange of $m$ and $n$, and it is zero for $m=n$. Therefore, we can combine the terms with $(n,m)$ and $(m,n)$ to rewrite
$\mathcal{S}^{(K)}[V]=\frac{2}{\hbar^{2}}\sum_{\begin{subarray}{c}nm\\ w_{m}^{(K)}\leq w_{n}^{(K)}\end{subarray}}(\epsilon_{m}-\epsilon_{n})^{2}\chi_{mn}^{(K)}w_{mn}^{(K)}\bigl|[V-\Phi]_{mn}\bigr|^{2}$
$\hskip 150.00023pt+\mathrm{const.},$ (56)
where we inserted $\chi_{mn}^{(K)}$ because if $m$ and $n$ satisfy $w_{m}^{(K)}=w_{n}^{(K)}$ and $m\neq n$, both $(m,n)$ and $(n,m)$ appear in the sum $\sum_{nm:w_{m}^{(K)}\leq w_{n}^{(K)}}$. The decomposition in Eq. (31) easily follows from Eq. (56).
Next, we show that the $K$-dependence of the partial action $\mathcal{T}^{(K,n)}[V]$ is between $O(K^{0})$ and $O(K^{-2})$. For simplicity, we neglect the $K$-dependence of $E^{(K)}$ because the dependence is subtle [see, e.g., Fig. 2(a)]. Then, the $K$-dependence of $\mathcal{T}^{(K,n)}[V]$ is due to the ratio $w_{mn}^{(K)}/w_{n}^{(K)}$. This ratio is written as
$\frac{w_{mn}^{(K)}}{w_{n}^{(K)}}=\frac{1}{K^{2}}\left[\sum_{s=0}^{K-1}\left(\frac{\epsilon_{m}-E^{(K)}}{\epsilon_{n}-E^{(K)}}\right)^{s}\right]^{2},$ (57)
where we used the second line of Eq. (48). The condition $w_{m}^{(K)}\leq w_{n}^{(K)}$ implies $|\epsilon_{m}-E^{(K)}|\leq|\epsilon_{n}-E^{(K)}|$, and thus, the sum in Eq. (57) is a finite geometric sum with a common ratio $(\epsilon_{m}-E^{(K)})/(\epsilon_{n}-E^{(K)})\in[-1,1]$. From Eq. (57), we deduce the following scaling of the ratio $w_{mn}^{(K)}/w_{n}^{(K)}$ in three extreme cases: (i) if $|\epsilon_{m}-E^{(K)}|\ll|\epsilon_{n}-E^{(K)}|$, the geometric sum converges quickly for a small value of $K$, and thus the ratio scales as $O(K^{-2})$; (ii) if $(\epsilon_{m}-E^{(K)})\simeq(\epsilon_{n}-E^{(K)})$ and $K$ is not too large, the geometric sum is roughly proportional to $K$, and the ratio scales as $O(K^{0})$; However, as $K$ becomes larger, the geometric sum starts to converge, and the scaling changes to $O(K^{-2})$; (iii) if $(\epsilon_{m}-E^{(K)})\simeq-(\epsilon_{n}-E^{(K)})$, the sign alternates in the geometric sum, and the value of $w_{mn}^{(K)}/w_{n}^{(K)}$ oscillates for a small $K$; Nevertheless, the scaling eventually becomes $O(K^{-2})$. A more general case lies between these extreme cases. The scaling of $\mathcal{T}^{(K,n)}[V]$ is determined by the sum of many terms with these scalings, and thus, the overall scaling also lies between $O(K^{0})$ and $O(K^{-2})$.
The asymptotic formula of $\mathcal{T}^{(K,n)}[V]$ is obtained from a similar consideration. Under the condition $w^{(K)}_{n}\geq w^{(K)}_{m}$, we can assume $(\epsilon_{n}-E^{(K)})^{K}\gg(\epsilon_{m}-E^{(K)})^{K}$ for a sufficiently large $K$. Using this relation in the second line of Eq. (22), we can approximate the ratio $w_{mn}^{(K)}/w_{n}^{(K)}$ as
$\frac{w_{mn}^{(K)}}{w_{n}^{(K)}}$ $\simeq\frac{1}{K^{2}(\epsilon_{n}-E^{(K)})^{2K-2}}\frac{(\epsilon_{n}-E^{(K)})^{2K}}{(\epsilon_{m}-\epsilon_{n})^{2}}$
$=\frac{1}{K^{2}}\frac{(\epsilon_{n}-E^{(K)})^{2}}{(\epsilon_{m}-\epsilon_{n})^{2}}.$ (58)
Inserting this approximation, the partial action $\mathcal{T}^{(K,1)}[V]$ is approximated by $K^{-2}(\epsilon_{1}-E^{(K)})^{2}\sum_{m:\,m\neq 1}|[V-\Phi]_{1m}|^{2}$, where we used the fact that $w_{1}^{(K)}$ becomes larger than any other $w_{m}^{(K)}$ for a sufficiently large $K$. This asymptotic formula of $\mathcal{T}^{(K,1)}[V]$ has been used in the main text to show that $\mathcal{S}^{(K)}[V]$ approaches the ideal action up to normalization.
Note that the ideal action can be derived more directly without relying on the partial action decomposition. In the limit $K\to\infty$, the fictitious Hamiltonian for the ground-state driving,
$\mathcal{P}^{\mathrm{GS},K}(H)=(H-E^{(K)})^{K}=\sum_{n=1}^{D}(\epsilon_{n}-E^{(K)})^{K}|\phi_{n}\rangle\langle\phi_{n}|,$ (59)
approaches $(\epsilon_{1}-E^{(K)})^{K}|\phi_{n}\rangle\langle\phi_{n}|$, which is proportional to the projection operator onto the ground state. Thus, $\mathcal{P}(\epsilon_{n})$ approaches $(\epsilon_{1}-E^{(K)})^{K}\delta_{n1}$. Inserting this into Eq. (14) reproduces the ideal action in Eq. (32) up to normalization and the constant term.
Appendix C Details of the computer algebra for general spin-1/2 systems